### v1

@@ 4.2.1 LLMs
  • The LLaMA family of models includes LLaMA-1 (Touvron et al., 2023a) and LLaMA-2 (Touvron et al., 2023b). We assess the performance of the “decapoda-llama-7b/13b-hf” and “meta-Llama-2-7b/13b-hf” models.
@@ 4.2.2 MLLMs
We evaluate a total of 23 Multimodal Large Language Models. The closed-source models include GPT-4V (OpenAI, 2023b), Gemini-V (Google, 2023), and Claude3 (Anthropic, 2024). The open-source MLLMs are developed by integrating pretrained visual models (PVMs) with LLM backbones. Typically, the PVMs are derived from pretrained CLIP models, such as ViT-L/14, ViT-g/14, etc. The LLMs span various families including the LLaMA family, with models like LLaMA-V1 and LLaMA-V2, Vicuna, the Flan-T5 family, and others.
@@
  • Fuyu-8B (AI, 2023) is a multimodal text and image converter trained by Adept AI [^8] https://www.adept.ai/. Architecturally, Fuyu is a pure decoder transformer - there is no image encoder and instead utilizes a transformer decoder. “adept/Fuyu-8B” is assessed.
@@
  • Qwen-VLs (Bai et al., 2023) are a series of highly performant and versatile vision-language foundation models based on Qwen-7B (Qwe, 2023) language model by introducing a new visual receptor. We evaluate the “Qwen-VL-Chat”.
@@ 4.3 Multimodal Datasets
Table 3: Evaluation datasets statistics for different tasks. “Labels” denotes the number of labels among each task. “-” means that the VQAMC task does not have a fixed label space, and each question has a unique answer.
@@
  • Visual Question Answering with Multimodal Contexts (VQAMC): ScienceQA (Lu et al., 2022a) is a well-known visual question-answering dataset featuring a variety of science topics. It provides both image-text contexts, which can range from semantically rich information to simple hints. Given our evaluation, we specifically utilize the portion of the dataset that includes image context. AlgoPuzzleVQA (Ghosal et al., 2024) is a novel dataset designed to challenge and evaluate the capabilities of MLLMs in solving algorithmic puzzles that require visual understanding, language comprehension, and complex algorithmic reasoning. The dataset includes areas such as boolean logic, combinatorics, graph theory, optimization, and search. It is crucial to note that the AlgoPuzzleVQA dataset requires the integration of textual context with visual elements to resolve specific queries effectively. Therefore, our assessment focuses exclusively on MLLMs capable of handling such intricate, dual-modality data interpretation. For simplicity, we refer to this dataset as “PuzzleVQA”. MMMU (Yue et al., 2023) presents significant challenges as an extensive benchmark designed for college-level multi-discipline multimodal understanding and reasoning. It encompasses six common disciplines: Art & Design, Business, Science, Health & Medicine, Humanities & Social Science, and Tech & Engineering. Our evaluations specifically focus on the validation set of MMMU, succinctly referred to as “MMMU-val.”
@@
  • Multimodal Aspect-Based Sentiment Analysis (MABSA): MABSA focuses on identifying sentiments associated with specific aspect terms within the given vision-text contexts (Hu et al., 2019; Yu and Jiang, 2019; Ju et al., 2021; Ling et al., 2022; Yang et al., 2022a; Yu et al., 2023a; Yang et al., 2022b; Yu et al., 2022). Our experiments utilize three widely recognized datasets: Twitter-2015 (Zhang et al., 2018), Twitter-2017 (Lu et al., 2018), and MASAD (Zhou et al., 2021). These datasets evaluate the effectiveness of MLLMs in accurately discerning sentiments related to specific aspects based on diverse multimodal contexts.
@@
Table 4: The best zero-shot performance, $A^{\tilde{i}}$ ($\uparrow$, measured by Accuracy), of various large language models on different datasets. Superscripts indicates the specific instruction that yields the best performance for each dataset on the respective model. The “Total” column represents the sum of accuracy scores across all datasets for each model. MVSA-S, MVSA-M, Twitter15, and Twitter17 refer to the MVSA-Single, MVSA-Multiple, Twitter-2015, and Twitter-2017 datasets, respectively. AdapterV2 and MultiGPT represent the LLaMA-AdapterV2 and Multimodal-GPT models, respectively. $\heartsuit$ indicates that the specific closed-source models (GPT-4V and Claude3-V) are evaluated on subsets. The highlighted text indicates the best performance on the full datasets. The highlighted text denotes superior performance on the subsets, outperforming the top results of the model evaluated across the full dataset. The notation “-” signifies that we do not assess LLMs for the PuzzleVQA and MMMU datasets. The exclusion is due to the inherent requirement of these two datasets for both image information and textual context to generate effective responses. The “Total” results of LLMs are also not calculated. The “Total” represents the aggregate results across all datasets, with the exception of PuzzleVQA and MMMU.

### v2

@@ 4.2.1 LLMs
  • The LLaMA family of models includes LLaMA-1 (Touvron et al., 2023a), LLaMA-2 (Touvron et al., 2023b), and LLaMA-3.1. We assess the performance of the “decapoda-llama-7b/13b-hf”, “meta-Llama-2-7b/13b-hf”, and “meta-llama-Llama-3.1-8B-Instruct” models.
@@ 4.2.2 MLLMs
  • GPT-4V (OpenAI, 2023b) extends ChatGPT’s capabilities to analyze image inputs, enhancing its functionality. We evaluate the “gpt-4-vision-preview” [^5] Our evaluation of GPT-4V via the corresponding API is carried out in February 2024. version.
  • GPT-4o (Hurst et al., 2024) is an autoregressive omni-model trained end-to-end across text, vision, and audio. This unified architecture processes all inputs and outputs through the same neural network, enabling seamless multimodal understanding. The “gpt-4o [^6] Our evaluation of GPT-4o via the corresponding API is carried out in February 2025.” is evaluated.
@@
The open-source MLLMs (31 models) are developed by integrating pretrained visual models (PVMs) with LLM backbones. Typically, the PVMs are derived from pretrained CLIP models, such as ViT-L/14, ViT-g/14, etc. The LLMs span various families including the LLaMA family, with models like LLaMA-V1 and LLaMA-V2, Vicuna, the Flan-T5 family, and others.
@@
  • Fuyu-8B (AI, 2023) is a multimodal text and image converter trained by Adept AI [^9] https://www.adept.ai/. Architecturally, Fuyu is a pure decoder transformer, which is no image encoder and instead utilizes a transformer decoder. “adept/Fuyu-8B” is assessed.
@@
  • LLaMA 3.2-Vision (Meta, 2024) comprises a collection of multimodal language models pretrained and instruction-tuned for visual tasks, including image recognition, reasoning, captioning, and visual question answering. In our evaluation, we focus on the “LLaMA-3.2-11B-Vision” model.
  • MiniCPM-V (Yao et al., 2024) is a series of efficient multimodal language models designed for edge devices, striking a balance between performance and computational efficiency. We evaluate “MiniCPM-V2.6”, the latest and most capable model in this series.
@@
  • GLM-4V (Hong et al., 2024) is a series of multimodal large language models (MLLMs) developed by Zhipu AI. These bilingual visual-language models are designed to explore image understanding capabilities in both English and Chinese. In our evaluation, we assess “THUDM/glm-4v-9b” and “glm-4v-plus”. Notably, glm-4v-plus [^10] Our evaluation of glm-4v-plus via the corresponding API is carried out in February 2025. is evaluated on a subset of data due to its high computational cost.
  • InternVL2.5-8B (Chen et al., 2024) retains the same model architecture as its predecessors, InternVL 1.5 and 2.0, following the “ViT-MLP-LLM” paradigm. The “OpenGVLab/InternVL2.5-8B” is assessed.
  • DeepSeek-VL2 (Wu et al., 2024b) is an advanced series of large Mixture-of-Experts (MoE) Vision-Language Models that significantly improves upon its predecessor. In this evaluation, the models “deepseek-ai/deepseek-vl2-tiny”, “deepseek-ai/deepseek-vl2-small”, and “deepseek-ai/deepseek-vl2” are assessed. Note that we use the API for the deepseek-vl2 model due to limited computing resources.
  • Qwen-VL Family (Bai et al., 2023) represents a series of highly performant and versatile vision-language foundation models derived from the Qwen series (Qwe, 2023). This family introduces new visual receptors, including Qwen-VL, Qwen2-VL (Wang et al., 2024), and the latest iteration, Qwen2.5-VL (Bai et al., 2025). In our evaluation, we focus on various model variants, i.e., “Qwen-VL-Chat” and “Qwen2-VL-7B-Instruct”, “Qwen2-VL-72B-Instruct”, “Qwen2.5-VL-3B-Instruct”, and “Qwen2.5-VL-7B-Instruct”. We also use the API for the Qwen2-VL-72B-Instruct model due to limited computing resources.
@@ 4.3 Multimodal Datasets
Table 3: Evaluation datasets statistics for different tasks. “Labels” denotes the number of labels among each task. “-” means that the VQAMC task has no fixed label space, and each question has a unique answer.
@@
  • Visual Question Answering with Multimodal Contexts (VQAMC): ScienceQA (Lu et al., 2022a) is a well-known visual question-answering dataset featuring a variety of science topics. It provides both image-text contexts, which can range from semantically rich information to simple hints. Given our evaluation, we specifically utilize the portion of the dataset that includes image context. AlgoPuzzleVQA (Chia et al., 2024) is a novel dataset designed to challenge and evaluate the capabilities of MLLMs in solving algorithmic puzzles that require visual understanding, language comprehension, and complex algorithmic reasoning. The dataset includes areas such as boolean logic, combinatorics, graph theory, optimization, and search. It is crucial to note that the AlgoPuzzleVQA dataset requires the integration of textual context with visual elements to resolve specific queries effectively. Therefore, our assessment focuses exclusively on MLLMs capable of handling such intricate, dual-modality data interpretation. For simplicity, we refer to this dataset as “PuzzleVQA”. MMMU (Yue et al., 2023) presents significant challenges as an extensive benchmark designed for college-level multi-discipline multimodal understanding and reasoning. It encompasses six common disciplines: Art & Design, Business, Science, Health & Medicine, Humanities & Social Science, and Tech & Engineering. Our evaluations specifically focus on the validation set of MMMU, succinctly referred to as “MMMU-val”.
@@
  • Multimodal Aspect-Based Sentiment Analysis (MABSA): MABSA focuses on identifying sentiments associated with specific aspect terms within the given vision-text contexts (Hu et al., 2019; Yu & Jiang, 2019; Ju et al., 2021; Ling et al., 2022; Yang et al., 2022a; Yu et al., 2023a; Yang et al., 2022b; Yu et al., 2022). Our experiments utilize three widely recognized datasets: Twitter-2015 (Zhang et al., 2018), Twitter-2017 (Lu et al., 2018), and MASAD (Zhou et al., 2021). These datasets evaluate the effectiveness of MLLMs in accurately discerning sentiments related to specific aspects based on diverse multimodal contexts.
@@
5 Experimental Results and Analysis
We evaluate the zero-shot performance of 45 models (as detailed in Table 2) across 16 datasets (listed in Table 3), utilizing 10 distinct instructions per model (as shown in Figures 3, 4, and 5). The related results in a total of 7200 (45 $\times$ 16 $\times$ 10) experimental results, each yielding an accuracy score denoted as $acc_{mdi}$ from Eq. 1. Although most of the evaluated datasets could rely on text context for inference, PuzzleVQA, and MMMU are exceptions, necessitating multimodal contexts. For ScienceQA, which primarily depends on the image modality to answer questions, we also explore the potential to extract knowledge from models using solely the image modality. Our comprehensive evaluation incorporates a range of metrics, including the Best Performance metric, the Mean Relative Gain metric, the Stability metric, and the Adaptability metric. Detailed analyses of these metrics and insightful conclusions related to them are presented in the subsequent sections.
5.1 Best Performance
Table 4: The best zero-shot performance, $A^{\tilde{i}}$ ($\uparrow$, measured by Accuracy), of various large language models on different datasets. Superscripts indicate the specific instruction that yields the best performance for each dataset on the respective model. The “Total” column represents the sum of accuracy scores across all datasets for each model. MVSA-S, MVSA-M, Twitter15, and Twitter17 refer to the MVSA-Single, MVSA-Multiple, Twitter-2015, and Twitter-2017 datasets, respectively. AdapterV2 and MultiGPT represent the LLaMA-AdapterV2 and Multimodal-GPT models, respectively. $\heartsuit$ indicates that the specific closed-source models (GPT-4V and Claude3-V) are evaluated on subsets. The highlighted text indicates the best performance on the full datasets. The highlighted text denotes superior performance on the subsets, outperforming the top results of the model evaluated across the full dataset. The notation “-” indicates datasets (PuzzleVQA and MMMU) where LLM assessment was not conducted, as these datasets inherently require visual and textual inputs for meaningful response generation. The “Total” denotes the aggregated performance across all evaluated datasets, while “Total” represents the aggregate results, excluding the VQAMC task (ScienceQA, PuzzleVQA, and MMMU).